How many liters of a 40% acid is needed to combine with a 15% solution to make 30 liters of a 20% solution?

Answer by checkley77(12844)

.40x+.15(30-x)=.20*30
.40x+4.5-.15x=6
.25x=6-4.5
.25x=1.5
x=1.5/.25
x=6 liters of 40% acid is used.
30-6=24 liters of 15% acid is used.
Proof:
.40*6+.15*24=.2*30
2.4+3.6=6
6=6